OpenStudy (anonymous):

What is the solution to p over zero point three + 1 = 8? 23.3 30 2.1 2.7

OpenStudy (anonymous):

\[\frac{ p }{0.3 }+1=8,\frac{ p }{ 0.3 }=8-1=7,p=7*0.3=?\]

OpenStudy (anonymous):

so 2.1

OpenStudy (anonymous):

correct.